OpenStudy (anonymous):

7-square root of 5 / 7+square root of 5

OpenStudy (a_clan):

multiply and divide by conjugates

OpenStudy (anonymous):

can u walk me trhoguh it

OpenStudy (a_clan):

yes

OpenStudy (a_clan):

7-(5)½ 7-(5)½ ------ * ------- [multiply both numerator and denominator 7+(5)½ 7-(5)½ by the conjugate of denominator]

OpenStudy (a_clan):

conjugate of (a + b) is (a-b)

OpenStudy (a_clan):

now in denominator use formula (a+b)(a-b)= (a²-b²)

OpenStudy (a_clan):

and in numerator use formula (a-b)(a-b)=(a-b)²

OpenStudy (a_clan):

then simplify

OpenStudy (a_clan):

This is the standard procedure for solving such problems